OpenStudy (anonymous):

Find the slope of the line passing through the points (-12, 1)(-6, 7).

OpenStudy (anonymous):

@Venny

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

Use the Slope Y2-Y1/X2-X1

OpenStudy (aum):

Slope = \(\Large \frac{y_2-y_1}{x_2-x_1} = \frac{7-1}{-6-(-12)} = ?\)

OpenStudy (anonymous):

6/6?

OpenStudy (aum):

6/6 = 1. Slope = 1.

OpenStudy (anonymous):

thanks

OpenStudy (aum):

yw.